What part of the job can Hanz do per hour if the whole job takes.
1. 5 hours
2. x hours
3. (x + 1) hours
4. 1/2 hour
5. 2/3 hour

Answer by math_helper(2461)


If he does the job in 1 hr he can do 1/1 = 1 job per hour
If he does the job in 2 hrs he can do 1/2 = 1/2 job per hour
--
The pattern is if x hours to do the job, he does 1/x job per hour.
--
1. If he takes 5 hours, he can do 1/5 job per hour
--
I'll let you finish, it will be more rewarding for you if you finish it.